# VEGGIE [Uncaptioned image]: Instructional Editing and Reasoning of Video Concepts with Grounded Generation

### v2

@@ 4 Experiments
[…]
@@
4.1 VEG-Bench and Metrics
As no existing benchmark is designed for fine-grained instructional video editing skills, we manually collect and annotate VEG-Bench, containing 132 video-instruction pairs that balanced cover 8 different video generative skills (15--20 for each). Beyond standard metrics, including text-to-video alignment (CLIP-Text [54]), video smoothness (CLIP-F [54]), and image quality (MUSIQ [32]), we also first introduce MLLM-as-a-Judge to give a holistic evaluation score according to the given original video, edited video, and user instruction. It is achieved by prompting GPT-4o [21] to evaluate whether the requested semantic change has been fulfilled, using a scale from 1 to 10. For addition and removal, we also introduce an object detector (GroundingDiNo [44]) to detect if the object is added/removed faithfully. For grounding and reasoning segmentation, we following video grounding tasks [12, 33, 57, 14] and adopt the Jaccard index ($\mathcal{J}$[31], F-measure ($\mathcal{F}$[13], and their mean ($\mathcal{J\&F}$). We also compute SSIM between the generated video and the original video masked with GT masks. More evaluation/metrics details are included in Appendix.
4.2 Experimental Setup
Settings Removal (FVD $\downarrow$) Grounding (SSIM $\uparrow$)
Grd.-only - 52.34
Rmv.-only 1098.52 -
Mixed 987.80 55.21
Figure 4: Visualization comparison between single-skill models and multi-skill models with mixed data training. We find tasks like grounding and removal can help each other.
Implementation Details. Our MLLM is initialized with LLaVA-OneVision-7B (LLaVA-OV) [38]. Our VidDM is initialized from the instructional image editing model, MagicBrush [82] with Stable Diffusion v1.5 backbone [56]. We further inflated 2D convolution layers with temporal attention layers, following AnimateDiff [22] to adapt videos. Our VEGGIE adopts a 2-stage curriculum training strategy (Sec. 3.2). In the first stage, we fully fine-tune the 2D convolution layers in the UNet, the alignment network, and the task query tokens in the MLLM on image data, with 862M trainable parameters. In the second stage, we train all 3 dimensions in the UNet, the alignment network, the task query tokens, and a LoRA in the MLLM, leading to 1.3B trainable parameters. Both stages are trained end-to-end with only a diffusion loss. More details are in the Appendix.
Baselines. We primarily compare our model with strong instructional editing models [19, 9, 65]. Additionally, we include non-instructional editing models [40, 10, 20] for completeness, although these are not fair baselines since they are not end-to-end and rely on additional conditions, such as depth maps or intermediate captions.
4.3 Experimental Results
Instructional Video Editing over Diverse Skills. As shown in Tab. 2, we evaluate 7 different models on VEG-Bench across 8 distinct editing skills. Overall, VEGGIE demonstrates the best performance among instructional video editing models. Compared to VEGGIE, non-instructional models often struggle with concept removal and addition. This limitation arises because these models rely on attention control or additional conditions (e.g., depth maps) that impose strong priors, constraining the model and making object addition or removal challenging. We also observe that InsV2V achieves high scores in quality and smoothness metrics, but underperforms in alignment and MLLM judgment, which demand faithful semantic changes. Qualitative examples in Fig. 3 illustrate that InsV2V often makes minimal changes to the input video, resulting in high video quality but unfaithful outputs. In contrast, VEGGIE strikes a better balance, delivering both high-quality visuals and accurate semantic alignment with the intended edits.
@@ 4.3 Experimental Results
Video Concept Grounding & Reasoning Segmentation We include additional results on video concept grounding and reasoning segmentation in Tab. 3. VEGGIE outperforms the diffusion-based baseline by a significant margin, showcasing its superior ability to accurately locate fine-grained object references and handle complex reasoning tasks. We hypothesize that through grounded generation, VEGGIE demonstrates remarkable precision in concept editing. For example, as shown in Fig. 10 in the Appendix, VEGGIE can remove the woman without altering the nearby girl.
4.4 Further Analyses and Discussions
Can Multi-Task Help Each Other? To test the previous hypothesis, we train our model on the VPLM [74] dataset, which includes paired grounding and removal tasks (approximately 5.5K samples for each task). We focus on these tasks as representative examples due to their straightforward evaluation against ground truth. As shown in Table 4, multitask training yields a lower FVD score and a higher SSIM score, demonstrating that learning to locate and remove a video concept can mutually reinforce performance. We show an example in Fig. 4. However, this conclusion only holds with a balanced data combination. We also observe that an excessive amount of grounding data can introduce more artifacts and negatively impact visual editing skills.
Task Query Visualization & Analysis via t-SNE. To analyze task/skill correlations, we project their grounded queries into lower-dimensional spaces using PCA and t-SNE. As shown in Fig. 5, distinct clusters form for each category (e.g., Addition), indicating effective differentiation by the model. Reasoning and Grounding appear together on the right. It may be because they both require cognitive/semantic understanding or logical reference. Color, Env, and Change clusters are closer to each other, indicating that the model views them as similar operations focusing on changing different visual attributes. Style lies in the lower-left region but remains relatively close to Color, Env, and Change. This proximity may reflect that “stylization” is conceptually similar to these visual attribute tasks, although it targets different transformations. Removal stands apart on the top, especially distant from Addition, indicating the model perceives them as distinct rather than inverse operations. In contrast, Addition lies closer to tasks like Reasoning and Grounding. It suggests that the act of adding elements may rely on similar semantic or referential processes (e.g., deciding what to add and how to reference the newly added element).
5 Conclusion
We present VEGGIE, a unified end-to-end model for instructional video editing that handles diverse pixel-level tasks, including concept addition, removal, style transfer, grounding, and reasoning segmentation. VEGGIE leverages MLLM for robust instruction understanding and employs a video diffusion model to execute pixel-level edits. Our framework uses a single diffusion loss for end-to-end optimization across varied tasks/skills unified in a consistent format. We also introduce a novel synthetic data generation pipeline and VEG-Bench, a benchmark that assesses a broad range of editing skills. Our VEGGIE outperforms previous state-of-the-art methods as a versatile, all-in-one solution. We hope our model, data, and benchmark to advance research on instructional generative video models.
Supplementary Material
6 Appendix
In this Appedix, we provide extra details on
  • Limitation and future work discussion.
6.1 Implementation Details
Training Details. We keep the VAE encoder and decoder frozen during the entire training process. In the first stage, we keep the MLLM (including visual encoder, MLP projector, and LLM) frozen, and fully fine-tune learnable grounded task queries, alignment network, and diffusion model, leading to around 800M training parameters. We set $1e^{-4}$ learning rate, and 96 batch size on each GPU. We use 16 A100 GPUs for the first stage of fine-tuning with 25K steps. In the second stage, we insert LoRA [27] modules into the LLM backbone, and inflate diffusion models by inserting extra temporal layers as in AnimateDiff [22]. We fine-tune LoRA, alignment network, learnable grounded task query tokens, and the diffusion model, leading to around 1.3B trainable parameters. We set $5e^{-4}$ learning rate, and 1 batch size with 8 gradient accumulation steps on 32 A100 GPUs. For LoRA, we set lora rank 64, lora alpha 16, and lora dropout 0.05. We train the second stage video model 2.5K step with 8 uniformly sampled frames.
Evaluation and Baseline Details. We randomly sample 3 seeds for both our method and baseline methods. In our experiments, we use different classifier-free guidance scores ($g_{T}$ and $g_{V}$ in Sec. 3.2) for different skills. Specifically, we set $g_{T}=14.5$ and $g_{V}=1.5$ for grounding and reasoning segmentation, while for other editing skills, we use $g_{T}=10.5$ and $g_{V}=2.0$.
6.2 Data Collection Details
As mentioned in the earlier Sec. 3.3, beyond collecting existing data, we proposed a novel data synthesis pipeline to generate instructional video data by animating images in the instructional image dataset.
We first use QWen2-VL [60] to caption the original image and give an animation prompt to animate the image via CogVideX1.5-I2V [73]. Please refer Tab. 5 and Tab. 6 to our prompt for caption and animation. After getting the animated video, we utilize AnyV2V [35] to edit the video based on the reference image (edited image from image dataset). The reference image gives a strong prior to maintaining the image dataset’s high-quality edit and thus transfer it to the video via the video editing model.
Next, we filter out videos by evaluating VBench metrics [30], including aesthetic quality, motion smoothness, image quality, subject consistency, and background consistency. We set thresholds at 0.6 for aesthetic quality, 65 for imaging quality, 0.9 for motion smoothness, subject consistency, and background consistency. We provide our generated data visualization in Fig. 8.
@@ 6.3 Limitation and Future Works
User You are an evaluator for instructional video editing tasks. Your job is to assess how well the edited video fulfills the user’s specific instructions. I will provide: 1. The original video (first GIF) 2. The edited video (second GIF) 3. The user’s instruction: [user instruction] Please evaluate the editing result using the following format: INSTRUCTION: [Repeat the user’s instruction] EVALUATION: - Accuracy score (1-10): [Your score] - Quality score (1-10): [Your score] - Appropriateness score (1-10): [Your score] - Overall score (1-10): [Your final score] EXPLANATION: [Provide a brief justification for your scores, highlighting specific strengths and weaknesses of the edit] RECOMMENDATION: [Optional suggestions for improvement] When scoring, consider: - Accuracy: Does the edit precisely follow the given instruction? - Quality: Is the edit visually seamless and natural-looking? - Appropriateness: Does the edit maintain coherence with the original video context? The overall scale is: 1-3: Poor - Major issues with the edit 4-6: Acceptable - Follows instruction but with noticeable flaws 7-8: Good - Clear, effective edit with minor issues 9-10: Excellent - Flawless execution of the instruction   Assistant Scores, Explanation, Recommendation
@@ 6.4 Extra Visualization
We provide extra visualization in Figs. 10, 9, 11, 12, 13, 14 and 15
Figure 9: Concept Addition.
Figure 10: Concept Removal.
Figure 11: Object Changes.
Figure 12: Style.
Figure 13: Environment and Background.
Figure 14: Visual Features.
Figure 15: Grounding.
Figure 16: Reasoning Segmentation.

### v3

@@ 6.3 More Quantative Results & Discussion
$\mathcal{J}$ $\mathcal{F}$ $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J\&F}$
@@
Video Concept Grounding & Reasoning Segmentation We include additional results on video concept grounding and reasoning segmentation in Tab.˜4. VEGGIE outperforms the diffusion-based baseline by a significant margin, showcasing its superior ability to accurately locate fine-grained object references and handle complex reasoning tasks. We hypothesize that through grounded generation, VEGGIE demonstrates remarkable precision in concept editing. For example, as shown in Fig. 10 in the Appendix, VEGGIE can remove the woman without altering the nearby girl.
@@ 6.4 Limitation and Future Works
Task Query Visualization & Analysis via t-SNE. To analyze task/skill correlations, we project their grounded queries into lower-dimensional spaces using PCA and t-SNE. As shown in Fig.˜7, distinct clusters form for each category (e.g., Addition), indicating effective differentiation by the model. Reasoning and Grounding appear together on the right. It may be because they both require cognitive/semantic understanding or logical reference. Color, Env, and Change clusters are closer to each other, indicating that the model views them as similar operations focusing on changing different visual attributes. Style lies in the lower-left region but remains relatively close to Color, Env, and Change. This proximity may reflect that “stylization” is conceptually similar to these visual attribute tasks, although it targets different transformations. Removal stands apart on the top, especially distant from Addition, indicating the model perceives them as distinct rather than inverse operations. In contrast, Addition lies closer to tasks like Reasoning and Grounding. It suggests that the act of adding elements may rely on similar semantic or referential processes (e.g., deciding what to add and how to reference the newly added element).
@@
User You are an evaluator for instructional video editing tasks. Your job is to assess how well the edited video fulfills the user’s specific instructions. I will provide: 1. The original video (first GIF) 2. The edited video (second GIF) 3. The user’s instruction: [user instruction] Please evaluate the editing result using the following format: INSTRUCTION: [Repeat the user’s instruction] EVALUATION: - Accuracy score (1-10): [Your score]
- Quality score (1-10): [Your score]
- Appropriateness score (1-10): [Your score] - Overall score (1-10): [Your final score] EXPLANATION: [Provide a brief justification for your scores, highlighting specific strengths and weaknesses of the edit]
RECOMMENDATION: [Optional suggestions for improvement]
When scoring, consider: - Accuracy: Does the edit precisely follow the given instruction? - Quality: Is the edit visually seamless and natural-looking? - Appropriateness: Does the edit maintain coherence with the original video context? The overall scale is: 1-3: Poor - Major issues with the edit 4-6: Acceptable - Follows instruction but with noticeable flaws 7-8: Good - Clear, effective edit with minor issues 9-10: Excellent - Flawless execution of the instruction  Assistant Scores, Explanation, Recommendation
@@ 6.5 Extra Visualization
We provide extra visualization in Figs.˜10, 9, 11, 12, 13, 14 and 15
Figure 9: More Examples of Concept Addition.
Figure 10: More Examples of Concept Removal.
Figure 11: More Examples of Object Changes.
Figure 12: More Examples of Stylization.
Figure 13: More Examples of Environment and Background Editing.
Figure 14: More Examples of Visual Features Editing.
Figure 15: More Examples of Object Grounding.
Figure 16: More Examples of Object Reasoning Segmentation.